OpenStudy (anonymous):

Simplify the expression. (–5 – c)(–1)

OpenStudy (katrinakaif):

First distribute -1 to -5 and -c

OpenStudy (anonymous):

(–5 – c)(–1) =(-5)(-1)-c(-1) =5+c